an 18 year old runner can complete a course in 13.8km with an average speed of 4.58m/s a 50 year old runner can cover the same distance with an average speed of 3.97m/s how much later in seconds should the younger runner start in order to finish the course at the same time as the older runner
1
Expert's answer
2020-04-02T11:00:55-0400
"t=t_2-t_1=\frac{s}{v_2}-\frac{s}{v_1}"

"t=\frac{13800}{3.97}-\frac{13800}{4.58}=463\ s"